I am using the formula w = Cr^-2. The value for w is 100 and the value for r is 3,963. In solving for C, I get a very large number.... 1,570,536,900. Is this correct?
Answer by checkley77(12844)

w = Cr^-2
C=W/R^-2
C=WR^2
C=100*3963^2
C=100*15,705,369
C=1.570,536,900 ANS.